A project manager is concerned how long her project will take so she has performed a Monte Carlo simulation. The project duration is affected to a large extent by whether or not it will rain during the project. If it does not rain then the project is fairly predictable. However, if it does rain then the project duration will be strongly affected by when and how much it rains. She calculates 10 realisations of the project assuming that it rains obtaining durations of 18, 18, 37, 26, 21, 36, 22, 27, 27 and 40, and 5 realisations assuming that it does not rain obtaining 13, 14, 14, 14 and 18, There is a 34% chance that it will rain Based on these results what is the expected duration of the project? (Answer to 2 decimal places) Number
Answer #1

Answer to the question)

Average projection if it rains = (18+18+37+26+21+36+22+27+27+40)/10

Average projection (rain) = 27.2

Probability of rain P(rain) = 0.34

.

Average projection when it doesnot rain = (13+14+14+14+18)/5

Average Projection ( no rain) = 14.6

Probability of no rain = P(no rain) = 1 - 0.34 = 0.66

.

Expected value = sum of product of average projection and respective probabilities

Expected value = 27.2*0.34 + 14.6 *0.66

Expected value = 18.88